# Weight-Entanglement Meets Gradient-Based Neural Architecture Search

### v1

@@ 6.0.1 Acknowledgments
This research was partially supported by the following sources: TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215; the German Federal Ministry of Education and Research (BMBF, grant RenormalizedFlows 01IS19077C); the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828; the European Research Council (ERC) Consolidator Grant “Deep Learning 2.0” (grant no. 101045765). Robert Bosch GmbH is acknowledged for financial support. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the ERC. Neither the European Union nor the ERC can be held responsible for them.
Appendix A Training across data fractions
Search Type Optimizer Train portion CIFAR10 CIFAR100
Single-Stage TangleNAS 50% $97.715\pm 0.0877$ $82.5375\pm 0.11764$
80% $\mathbf{97.872\pm 0.05369}$ $\mathbf{82.6675\pm 0.16073}$
Two-Stage SPOS+RS 50% $97.6799\pm 0.0262$ $82.5372\pm 0.27954$
80% $97.7674\pm 0.0239278$ $82.210\pm 0.14242$
SPOS+RE 50% $97.77\pm 0.03752$ $82.354\pm 0.11974$
80% $97.6425\pm 0.02286$ $82.5175\pm 0.1135$
Table 9: Evaluation on the AutoFormer-T space for CIFAR10 and CIFAR100
@@ Appendix A Training across data fractions
CIFAR10 CIFAR100 ImageNet16-120
Single-Stage DrNAS 50% WS $\mathbf{94.36\pm 0.00}$ $72.245\pm 0.7315$ $\mathbf{46.37\pm 0.00}$
80% $\mathbf{94.36\pm 0.00}$ $71.1525\pm 0.697$ $\mathbf{46.37\pm 0.00}$
@@
Two-Stage SPOS+RS 50% WE $89.107\pm 0.88392$ $56.865\pm 2.5968$ $31.665\pm 1.1460$
80% $87.778\pm 2.4462$ $53.68\pm 4.1736$ $30.5449\pm 3.6425$
SPOS+ES 50% WE $87.133\pm 2.6050$ $56.463\pm 2.3417$ $29.7849\pm 3.0149$
80% $89.095\pm 0.8250$ $56.363\pm 4.7244$ $30.935\pm 3.54648$
Table 10: Comparison of test accuracies of single and two stage methods with WS and WE on NB201 search space
Table 11: Evaluation on toy cell-based search space on fashion-mnist dataset
Search Type Optimizer Train portion Supernet type Test acc
Single-Stage DrNAS 50% WS $91.19\pm 0.0490$
80% $91.125\pm 0.0334$
@@
80% $91.065\pm 0.1633$
Two-Stage SPOS+RS 50% WE $90.68\pm 0.25345$
80% $90.687\pm 0.10969$
SPOS+ES 50% WE $90.3175\pm 0.2233$
80% $90.595\pm 0.2188$
@@
Search Type Optimizer Train portion Supernet type Test acc
Single-Stage DrNAS 50% WS 10%
80% 10%
@@
80% 82.495±.0.46074plus-or-minus82.495.0.4607482.495$\pm$.0.4607482.495 ± .0.46074
Two-Stage SPOS+RS 50% WE $81.253\pm 0.67174$
80% $81.345\pm 0.38336$
SPOS+ES 50% WE $81.890\pm 0.8002$
80% $82.322\pm 0.60356$
@@
Table 11: Evaluation on toy cell-based search space on fashion-mnist dataset
Table 12: Evaluation on toy conv-macro search space on CIFAR10 dataset
Search Type Optimizer Train portion Supernet CIFAR10 ImageNet
Single-Stage TangleNAS 50% WE $\mathbf{2.556\pm 0.03406}$ $\mathbf{25.69}$
80% $2.67\pm 0.0756$ 25.742
DrNAS 50% WS $2.625\pm 0.0750$ 26.29
80% $2.580\pm 0.02798$ $\mathbf{25.67}$
Two-Stage SPOS+RS 50% WE $2.965\pm 0.07193$ 27.114
80% $2.965\pm 0.07193$ 27.114
SPOS+RE 50% WE $3.20000\pm 0.06557$ 27.424
80% $3.00249\pm 0.037052$ 26.76
Table 13: Comparison of test errors of single and two stage methods with WS and WE on DARTS search space
Appendix B Search and Training Times
B.1 NB201 and DARTS
Search Type Optimizer Supernet Search Time (hrs)
Single-Stage DrNAS WS 20.9166
TangleNAS WE 20.3611
Two-Stage SPOS+RS WE 29.51388
SPOS+ES WE 26.74721
Table 14: Comparison of search times NB201
Search Type Optimizer Supernet GPU MemSearch Time(hrs)
Single-Stage DrNAS WS 81.08% 29.4444
TangleNAS WE 52.26% 25.472
Two-Stage SPOS+RS WE 25.66% 18.7444
SPOS+ES WE 25.66% 14.794
Table 15: Comparison of test errors of single and two-stage methods on the DARTS search space
Appendix C Search Space details
@@
This particular search space takes its inspiration from the prominently used Differentiable Architecture Search (DARTS) space, and is composed of diminutive triangular cells, with each edge offering four choices of operations: (a) Separable 3x3 Convolution, (b) Separable 5x5 Convolution, (c) Dilated 3x3 Convolution, and (d) Dilated 5x5 Convolution. The macro-architecture of the model comprises of three cells of types reduction, normal, and reduction again stacked together. Notably, we entangle the 3x3 and 5x5 kernel weights for each operation type, i.e., separable convolutions and dilated convolutions. We evaluate these search spaces and their architectures on the Fashion-MNIST dataset by creating a small benchmark which we release here.
@@
This toy space draws its inspiration from MobileNet-like spaces where we search for the number of channels and the kernel size of convolutional layers in a network (also called macro search space) for four convolutional layers. Every convolutional layer has a choice of three kernel sizes and number of channels. See Table 16 for more details. We evaluate this search space on the CIFAR10 dataset by creating a small benchmark which we release here.
@@
Table 16: Toy Convolutional-Macro Search Space
NASBench201 Space
The NASBench201 [Dong and Yang, 2020] has a single cell type which is stacked 5 times. Each cell has 4 nodes and each node are connected by operations from amongst skip connection, conv 3x3, conv 1x1, max-pool or average-pool. Opportunities for weight entanglement in this space is limited and we entangle the conv 3x3 and conv 1x1 kernel on every edge for every cell, thus obtaining parameter savings over traditional weight sharing.
@@
The DARTS [Liu et al., 2019] search space has a cell with 13 possible edges stacked 8 times. The nodes of the cell are connected by operations amongst :none, max_pool_3x3, avg_pool_3x3, skip_connect, sep_conv_5x5, sep_conv_3x3, dil_conv_5x5, dil_conv_3x3.
@@
We present the details of the AutoFormer Space and the Language Model space is Table 19 and Table 19 respectively.
Table 17: Choices for AutoFormer-T Search Space
Table 18: Choices for Language Model Space
@@
Architectural Parameter Choices
Embedding Dimension 384, 576, 768
Number of Heads 6, 8, 12
MLP Expansion Ratio 2, 3, 4
Number of Layers 5,6,7
Architectural Parameter Choices
Kernel sizes (all layers) 3, 5, 7
Channel expansion (all layers) 3,4,6
Number of blocks 2,3,4
Table 17: Choices for AutoFormer-T Search Space
Table 18: Choices for Language Model Space
Table 19: MobileNet Search Space
Appendix D Details on Method
Traditional cell-based search spaces primarily consider independent operations (e.g., convolution or skip). One-shot differentiable optimizers thus have their mixture operations tailored to these search spaces which are not general enough to be applied to macro-level architectural parameters. Consider, e.g., the task of searching for the embedding dimension and the expansion ratio for a transformer. Here, a single operation, i.e., a linear expansion layer, has two different architectural parameters - one corresponding to the choice of embedding dimension and the other to the expansion ratio. To adapt single-stage methods to these combined operation choices in the search space, we propose the combi-superposition operation.
The combi-superposition operation simply takes the cross product of architectural parameters for the embedding dimension and expansion ratio and assigns its elements to every combination of these dimensions, allowing us to optimize jointly in this space without the need for a separate forward pass for each combination. Every combination maps to a unique sub-matrix of the operator weight matrix, indexed using both the embedding dimension and the expansion ratio. To address shape mismatches of the different operation weights during forward passes, every sub-matrix is zero-padded to match the shape of the largest matrix. See Algorithm 2 for more details, and Figure 2(b) for an overview of the idea.
@@ Appendix D Details on Method
  $\alpha,\beta\leftarrow\text{normalize}(\alpha),\text{normalize}(\beta)$Normalize architecture parameters (e.g., using softmax)}
  $\alpha\beta\leftarrow\alpha\times\beta$Cross product of $\alpha$ and $\beta$ for dimension $n\times m$}
  for $i\leftarrow 1$ to $n\times m$ do
     $W\_i=W[:(embed\_dim[i]\times expansion\_ratio[i]),:embed\_dim[i]]$
     $b\_i=b[:embed\_dim[i]\times expansion\_ratio[i]]$
     $W_{mix}=W_{mix}+\alpha\beta[i]\times\text{PAD}(W_{i})$
     $b_{mix}=b_{mix}+\alpha\beta[i]\times\text{PAD}(b_{i})$
@@
For completeness and for comparison with TangleNAS (Algorithm 1), we present Algorithm 3 which describes a generic two-stage method on a macro search space with WE. Also, vanilla single-stage methods on cell-based WS spaces follow Algorithm 4.
@@
2:  $Cell_{j}\leftarrow DAG(\mathcal{O_{j},\mathcal{W}_{\max_{j}}})$ /* defined for j=1...M */
3:  $Supernet\leftarrow\cup_{i}^{M}Cell_{i}$
@@
12:   /* Search */ 13:  $Supernet^{*}\leftarrow$ pre-trained supernet 14:  $selected\_arch\leftarrow$ Evolutionary-Search($Supernet^{*}$)
@@
2:  $Cell_{i}\leftarrow DAG(\mathcal{O_{j},\mathcal{W_{O_{j}}}})$ /* defined for i=1...M */
3:  $Supernet\leftarrow\cup_{i}^{M}Cell_{i}\cup\mathcal{A}$
@@
Appendix E Experimental Setup
E.1 Toy Search Spaces

### v2

@@ Acknowledgments
This research was partially supported by the following sources: TAILOR, a project funded by EU Horizon 2020 research and innovation programme under GA No 952215; the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under grant number 417962828; the European Research Council (ERC) Consolidator Grant “Deep Learning 2.0” (grant no. 101045765). Robert Bosch GmbH is acknowledged for financial support. The authors acknowledge support from ELLIS and ELIZA. Funded by the European Union. Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the ERC. Neither the European Union nor the ERC can be held responsible for them.
Appendix A Submission Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? Yes
    2. (b)
      Did you describe the limitations of your work? Yes, in Appendix I
    3. (c)
      Did you discuss any potential negative societal impacts of your work? NA
    4. (d)
      Did you read the ethics review guidelines and ensure that your paper conforms to them? https://2022.automl.cc/ethics-accessibility/ Yes
  2. 2.
    If you ran experiments…
    1. (a)
      Did you use the same evaluation protocol for all methods being compared (e.g., same benchmarks, data (sub)sets, available resources)? Yes
    2. (b)
      Did you specify all the necessary details of your evaluation (e.g., data splits, pre-processing, search spaces, hyperparameter tuning)? Yes
    3. (c)
      Did you repeat your experiments (e.g., across multiple random seeds or splits) to account for the impact of randomness in your methods or data? Yes, for experiments on smaller scales
    4. (d)
      Did you report the uncertainty of your results (e.g., the variance across random seeds or splits)? Yes, for experiments on smaller scales
    5. (e)
      Did you report the statistical significance of your results? Yes, we perform a two tailed p-value test on the tinystories dataset for language modeling
    6. (f)
      Did you use tabular or surrogate benchmarks for in-depth evaluations? Yes, We used NAS-Bench-201 which is tabular
    7. (g)
      Did you compare performance over time and describe how you selected the maximum duration? Yes, we study performance over compute budgets
    8. (h)
      Did you include the total amount of compute and the type of resources used (e.g., type of gpus, internal cluster, or cloud provider)? Yes
    9. (i)
      Did you run ablation studies to assess the impact of different components of your approach? No, We only add a single main component in our approach
  3. 3.
    With respect to the code used to obtain your results…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results, including all requirements (e.g., requirements.txt with explicit versions), random seeds, an instructive README with installation, and execution commands (either in the supplemental material or as a url)? Yes
    2. (b)
      Did you include a minimal example to replicate results on a small subset of the experiments or on toy data? Yes
    3. (c)
      Did you ensure sufficient code quality and documentation so that someone else can execute and understand your code? Yes
    4. (d)
      Did you include the raw results of running your experiments with the given code, data, and instructions? Yes, we include the smaller pickle files for toy benchmarks we create and results logs wherever possible. We will share larger raw files eg: checkpoints upon acceptance, as anonymization of these resources is difficult.
    5. (e)
      Did you include the code, additional data, and instructions needed to generate the figures and tables in your paper based on the raw results? Yes
  4. 4.
    If you used existing assets (e.g., code, data, models)…
    1. (a)
      Did you cite the creators of used assets? Yes
    2. (b)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating if the license requires it? We do not use personal data
    3. (c)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? We do not use personal data
  5. 5.
    If you created/released new assets (e.g., code, data, models)…
    1. (a)
      Did you mention the license of the new assets (e.g., as part of your code submission)? Yes
    2. (b)
      Did you include the new assets either in the supplemental material or as a url (to, e.g., GitHub or Hugging Face)? Yes
  6. 6.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? NA
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (irb) approvals, if applicable? NA
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? NA
  7. 7.
    If you included theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? NA
    2. (b)
      Did you include complete proofs of all theoretical results? NA
Appendix B Training across data fractions
TangleNAS is more robust across dataset fractions for network weights and architecture optimization as seen from Tables 9, 10, 11, and 12.
Search Type Optimizer Train portion CIFAR-10 (%) CIFAR-100 (%)
Single-Stage TangleNAS 50% $97.715\pm 0.088$ $82.538\pm 0.118$
80% $\mathbf{97.872\pm 0.054}$ $\mathbf{82.668\pm 0.161}$
Two-Stage SPOS+RS 50% $97.680\pm 0.026$ $82.537\pm 0.280$
80% $97.767\pm 0.024$ $82.210\pm 0.142$
SPOS+ES 50% $97.77\pm 0.038$ $82.354\pm 0.120$
80% $97.642\pm 0.023$ $82.517\pm 0.114$
Table 9: Evaluation on the AutoFormer-T space for CIFAR-10 and CIFAR-100.
@@ Appendix B Training across data fractions
CIFAR-10 CIFAR-100 ImageNet16-120
Single-Stage DrNAS 50% WS $\mathbf{94.36\pm 0.000}$ $72.245\pm 0.732$ $\mathbf{46.37\pm 0.00}$
80% $\mathbf{94.36\pm 0.00}$ $71.153\pm 0.697$ $\mathbf{46.37\pm 0.00}$
@@
Two-Stage SPOS+RS 50% WE $89.107\pm 0.884$ $56.865\pm 2.597$ $31.665\pm 1.146$
80% $87.778\pm 2.446$ $53.68\pm 4.174$ $30.545\pm 3.643$
SPOS+ES 50% WE $87.133\pm 2.605$ $56.463\pm 2.342$ $29.785\pm 3.015$
80% $89.095\pm 0.825$ $56.363\pm 4.724$ $30.935\pm 3.546$
Table 10: Comparison of test accuracies of single and two-stage methods with WS and WE on NB201 search space.
Search Type Optimizer Train portion Supernet type Test acc (%)
Single-Stage DrNAS 50% WS $91.19\pm 0.049$
80% $91.125\pm 0.033$
@@
80% $91.065\pm 0.163$
Two-Stage SPOS+RS 50% WE $90.68\pm 0.253$
80% $90.687\pm 0.110$
SPOS+ES 50% WE $90.318\pm 0.223$
80% $90.595\pm 0.219$
@@
Table 11: Evaluation on toy cell-based search space on Fashion-MNIST dataset.
Search Type Optimizer Train portion Supernet type Test acc (%)
Single-Stage DrNAS 50% WS 10
80% 10
@@
80%   $82.495\pm.0.461$
Two-Stage SPOS+RS 50% WE $81.253\pm 0.672$
80% $81.345\pm 0.383$
SPOS+ES 50% WE $81.890\pm 0.800$
80%    $82.322\pm 0.604$
@@
Table 12: Evaluation on toy conv-macro search space on CIFAR-10 dataset.
Search Type Optimizer Train portion Supernet CIFAR-10 (%) ImageNet (%)
Single-Stage TangleNAS 50% WE $\mathbf{2.556\pm 0.034}$ $\mathbf{25.69}$
80% $2.67\pm 0.076$ 25.742
DrNAS 50% WS $2.625\pm 0.075$ 26.29
80% $2.580\pm 0.028$ $\mathbf{25.67}$
Two-Stage SPOS+RS 50% WE $2.965\pm 0.072$ 27.114
80% $2.965\pm 0.072$ 27.114
SPOS+ES 50% WE $3.200\pm 0.066$ 27.424
80% $3.002\pm 0.037$ 26.76
Table 13: Comparison of test errors of single and two-stage methods with WS and WE on DARTS search space.
Appendix C Comparison with different gradient-based optimizers
We also compare DrNAS against different gradient-based NAS optimizers in Tables 15, 14, and 16 on the toy cell-based, toy macro and the AutoFormer search spaces. We observe that DrNAS outperforms GDAS and DARTS on all of these search spaces, showing its robust nature.
Optimizer Test-acc (%)
TangleNAS+DrNAS 83.02
SPOS+RS 81.25
SPOS+ES 81.89
TangleNAS+DARTS_v1 81.61
TangleNAS+DARTS_v2 81.49
TangleNAS+GDAS 10 (degenerate)
Table 14: Comparison of DrNAS with other gradient-based optimizers on the toy conv-macro search space on CIFAR-10.
Optimizer Test-acc (%)
TangleNAS+DrNAS 90.930
SPOS+RS 90.688
SPOS+ES 90.595
TangleNAS+DARTS_v1 89.905
TangleNAS+DARTS_v2 90.747
TangleNAS+GDAS 90.618
Table 15: Comparison of DrNAS with other gradient-based optimizers on the toy cell-based search space on FashionMNIST dataset.
Optimizer CIFAR-10 CIFAR-100
TangleNAS+DrNAS 97.872 $\pm$ 0.054 82.668 $\pm$ 0.161
SPOS+RS 97.767 $\pm$ 0.024 82.210 $\pm$ 0.142
SPOS+ES 82.518 $\pm$ 0.114 82.518 $\pm$ 0.114
TangleNAS+DARTS_v1 97.672 $\pm$ 0.040 82.107 $\pm$ 0.392
TangleNAS+GDAS 97.45 $\pm$ 0.096 82.120 $\pm$ 0.281
Table 16: Comparison of DrNAS with other gradient-based optimizers on the AutoFormer-T space for CIFAR-10 and CIFAR-100.
Appendix D Search Space details
@@
This particular search space takes its inspiration from the prominently used Differentiable Architecture Search (DARTS) space, and is composed of diminutive triangular cells, with each edge offering four choices of operations: (a) Separable $\times3$ Convolution, (b) Separable $\times5$ Convolution, (c) Dilated $\times3$ Convolution, and (d) Dilated $\times5$ Convolution. The macro-architecture of the model comprises three cells of the types reduction, normal, and reduction again stacked together. Notably, we entangle the $\times3$ and $\times5$ kernel weights for each operation type, i.e., separable convolutions and dilated convolutions. We evaluate these search spaces and their architectures on the Fashion-MNIST dataset by creating a small benchmark, which we release here.
@@
This toy space draws its inspiration from MobileNet-like spaces where we search for the number of channels and the kernel size of convolutional layers in a network (also referred to as a macro search space) for four convolutional layers. Every convolutional layer has a choice of three kernel sizes and number of channels. See Table 17 for more details. We evaluate this search space on the CIFAR-10 dataset by creating a small benchmark, which we release here.
@@
Table 17: Toy Convolutional-Macro Search Space.
NAS-Bench-201 Space
The NAS-Bench-201 search space (dong2020bench) has a single cell type which is stacked 15 times, with residual blocks (he-cvpr16a) after the fifth and tenth cells. Each cell has 4 nodes, with each node connected to the previous ones by operations chosen from skip connection, $\times3$ convolution, $\times1$ convolution, $\times3$ average-pooling, and zeroize, which zeros out the input feature map. There is limited scope for weight-entanglement in this space, given that only two of the five candidate operations contain learnable parameters. In this case, we entangle the $\times3$ and $\times1$ convolution kernels on every edge for every cell, thus obtaining parameter savings over traditional weight sharing.
Architectural Parameter Choices
Embedding Dimension 384, 576, 768
Number of Heads 6, 8, 12
MLP Expansion Ratio 2, 3, 4
Number of Layers 5,6,7
Table 18: Choices for Language Model Space.
@@
The DARTS (liu-iclr19a) search space has two kinds of cells - normal and reduction. Reduction cells double the number of channels in its outputwhile halving the width and height of the feature maps. The supernet consists of 8 cells stacked sequentially, while the discretized network stacks 20 cells. Reduction cells are placed at 1/3 and 2/3 of the total depth of the network. Each cell takes two inputs - one each from the outputs of the two previous cells. There are 14 edges in each cell, with each edge in the supernet consisting of 8 candidate operations as follows: $\times3$ max pooling, $\times3$ average pooling, skip connect, $\times3$ separable convolutions, $\times5$ separable convolutions, $\times3$ dilated convolutions, $\times5$ dilated convolutions, and none (no operation).
Architectural Parameter Choices
Kernel sizes (all layers) 3, 5, 7
Channel expansion (all layers) 3,4,6
Number of blocks 2,3,4
Table 19: MobileNetV3 Search Space.
@@
We present the details of the AutoFormer Space and the Language Model space is Tables 20 and 18, respectively.
@@
Table 20: Choices for AutoFormer-T Search Space.
Appendix E Methodological details
Combi-Superposition
Traditional cell-based search spaces primarily consider independent operations (e.g., convolution or skip). One-shot differentiable optimizers thus have their mixture operations tailored to these search spaces, which are not general enough to be applied to macro-level architectural parameters. Consider, e.g., the task of searching for the embedding dimension and the expansion ratio for a transformer. Here, a single operation, i.e., a linear expansion layer, has two different architectural parameters – one corresponding to the choice of embedding dimension and the other to the expansion ratio. To adapt single-stage methods to these combined operation choices in the search space, we propose the combi-superposition operation.
The combi-superposition operation simply takes the cross product of architectural parameters for the embedding dimension and expansion ratio and assigns its elements to every combination of these dimensions. This allows us to optimize jointly in this space without the need for a separate forward pass for each combination. Every combination maps to a unique sub-matrix of the operator weight matrix, indexed by both the embedding dimension and the expansion ratio. To address shape mismatches of the different operation weights during forward passes, every sub-matrix is zero-padded to match the shape of the largest matrix. See Algorithm 2 for more details, and Figure 3 for an overview of the idea.
@@ Appendix E Methodological details
For ease of presentation, we show the algorithm for superimposing along two dimensions. However, in practice, we can superimpose an arbitrary number of dimensions (e.g., four dimensions in our experiments with NanoGPT).
@@
for $i\leftarrow 1$ to $n$ do
  for $j\leftarrow 1$ to $m$ do
   $W\_{ij}=W[:(embed\_dim[i]\times expansion\_ratio[j]),:embed\_dim[i]]$
   $b\_{ij}=b[:embed\_dim[i]\times expansion\_ratio[j]]$
   $W_{\text{mix}}\leftarrow W_{\text{mix}}+\text{normalize}(\alpha[i])\cdot\text{normalize}(\beta[j])\times\text{PAD}(W_{ij})$
   $b_{\text{mix}}\leftarrow b_{\text{mix}}+\text{normalize}(\alpha[i])\cdot\text{normalize}(\beta[j])\times\text{PAD}(b_{ij})$
  end for
@@
For completeness and to facilitate comparison with TangleNAS (Algorithm 1), we present Algorithm 3, which describes a generic two-stage method on a macro search space with weight entanglement (WE). Additionally, vanilla single-stage methods on cell-based weight sharing (WS) spaces follow Algorithm 4.
@@
[…]
@@
[…]
@@
[…]
@@
Compatibility issues between weight-entanglement and gradient-based methods
We address the incompatibility between gradient-based NAS methods and weight-entangled (WE) spaces as follows (where n refers to the number of operation choices):
  • Gradient-based NAS methods do not share or entangle weights among competing operations, which increases their GPU memory footprint. We tackle this issue by adopting weight-entanglement from two-stage methods, thereby reducing the parameter size of the supernet from $\mathcal{O}$}(n) to $\mathcal{O}$}(1).
  • Gradient-based NAS methods compute a weighted combination of output of the operations. Even after entangling the operation weights, this approach leads to increased GPU memory usage because all intermediate competing activations/features need to be preserved in memory. Additionally, this method does not scale well with an increase in the number of operation choices, as GPU memory consumption during forward propagation scales as $\mathcal{O}$}(n). To address this, we propose weight superposition, which computes the weighted combination directly within the space of entangled weights.
  • In vanilla gradient-based methods, a forward pass is computed independently for each operation choice, resulting in a time and memory complexity of $\mathcal{O}$}(n). In contrast, TangleNAS computes only a single forward pass using the superimposed weights, reducing the cost of the forward pass to $\mathcal{O}$}(1).
Details on Figure 1
In the overview Figure 1, we use rounded squares to represent the input and output feature maps of a convolution. The colored cubes represent the convolutional kernels, while the colored rectangles denotes non-convolutional architectural choices. These two differ primarily in how their weights are entangled.
Consider Figure 1 (a), which illustrates the forward pass of an input feature map through the candidate operations on an edge of the supernet in a two-stage weight-entanglement method (such as OFA). In two-stage methods, random paths through the supernet are sampled and trained in the first stage. In this figure, thick lines indicate the paths sampled in a given step. The weights of the operation choices, depicted in the figure by colored cubes and rectangles, overlap with one another, showing that all the operations use slices of a common, large weight matrix. We show three choices of convolutions, each of a different kernel size (say, $\times1$, $\times3$, and $\times5$), represented by cubes of varying sizes. The $\times1$ and $\times3$ convolutions use slices of the larger $\times5$ convolution as their weights, represented in the figure by nesting the smaller kernels inside the larger ones. Since only one operation is sampled at a time (in this case, the orange kernel), there is only one output feature map. This feature map, after global average pooling, is then passed through one of the three non-convolutional operation choices. The operation choices here may be the embedding dimension, for example, and different slices of the largest feedforward network are used for the choices of embedding dimensions. At the end of the first stage in Figure 1 (a), the supernet has been trained along different paths. In the second stage, paths are sampled from the trained supernet using black-box methods (such as random or evolutionary search) and evaluated on the dataset to obtain the optimal architecture, which we represent with optim arch.
Now, consider Figure 1 (b), which represents the forward pass in a single-stage weight-sharing method (such as DARTS). As you can see, there is no nesting of the weights of the three convolutions or feedforward networks in this case, indicating that each has its own distinct set of weights. Naturally, this will incur more GPU memory usage, as more weights need to be stored and their gradients computed. The outputs from each of these convolutional (or feedforward) operations are then weighted by $\alpha_{1}$, $\alpha_{2}$, and $\alpha_{3}$ (or $\beta_{1}$, $\beta_{2}$, and $\beta_{3}$), which represent the architectural parameters of the operations. These weighted feature maps are then summed up to produce the output feature maps.
The optimization loop of single-stage weight-sharing methods, shown in Figure 1(b), aims to find the optimal values for the architectural parameters $\alpha$ and $\beta$. We obtain the discretized model by selecting only the operations in the supernet with the highest values of these parameters at the end of this loop. Specifically, the convolutional kernel and non-convolutional operator (in this case, the feedforward network) with the maximum values of $\alpha$ and $\beta$ are represented as argmax($\alpha_{1}$, $\alpha_{2}$, $\alpha_{3}$) and argmax($\beta_{1}$, $\beta_{2}$, $\beta_{3}$), respectively. The resulting discretized model, or optimal architecture, is denoted as optimal arch. Note that these methods do not require black-box search, as the optimal architecture can be obtained directly from the learned architectural parameters.
In Figure 1 (c), we depict our hybrid framework, which utilizes both entangled (nested) weights and architectural parameters. This approach reduces GPU memory usage due to weight entanglement and allows for the optimal architecture to be obtained directly from the learned parameters, eliminating the need for a black-box search.
Appendix F Experimental Setup
F.1 Toy Search Spaces
Below are the hyperparameter settings for the two toy spaces for TangleNAS and SPOS. All experiments were run on a single RTX-2080 GPU.